# Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late EoR

### v1

@@ 2.3.1 Generating density and collapsed halo field
For generating the cosmological density field, we use the approach adopted in the publicly available semi numerical reionization model 21cmFAST (Mesinger et al., 2011). The initial conditions on density and velocity fields are generated in Lagrangian space following the gravitational evolution of the fields via Zel’Dovich Approximation (Zel’dovich, 1970). This approximation has been proven to be successful in closely resembling with the full N body simulations with less computational cost. We use a simulation of box size $256\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$ which is sufficiently large to serve the purpose of this study. After having the particle fields, we gridded those at a high resolution ($0.5h^{-1}\mathrm{cMpc}$) and then averaged it to $\Delta x=4h^{-1}\mathrm{cMpc}$ which is the resolution we adopt throughout the study.
@@
Besides, to construct the collapsed halo mass fraction, we adopt the Excursion Set Formalism (ESF-L) in Lagrangian space, following another semi numerical reionization model AMBER (Trac et al., 2022). To be specific, the formalism uses Extended Press Schecter theory which computes the collapsed mass fraction ($f_{\mathrm{coll}}$) above a minimum threshold halo mass ($M_{\mathrm{min}}$) given the filtered overdensity at a cosmological scale. Here, we use a spherical tophat filter for filtering the overdensity at a certain scale ($2\leavevmode\nobreak\ \mathrm{cMpc}$), which provides a good match with the full N body simulation. At each redshift, we generated collapsed fraction field for a logarithmic minimum mass ($\log M_{\mathrm{min}}$) interval of $0.1$ in the range $7-12$.
@@
In the left panel of Figure 4, we show the overdensity distribution of Nyx simulation (blue) and the generated overdensities using semi numerical method (dashed orange) as discussed above. We checked that the nature of these two distributions are in agreement with each other. This resemblance is useful to avoid any discrepancies in optical depth due to the background density field.
2.3.2 Semi numerical reionization Model
@@
We utilize an explicitly photon conserving semi numerical reionization model, SCRIPT which provides the ionization and thermal state of the universe in a cosmologically representative simulation volume (Choudhury & Paranjape, 2018; Maity & Choudhury, 2022a). To initiate the model, we provide inputs comprising the density field and the distribution of collapsed halos capable of emitting ionizing radiation as generated by the method discussed earlier. For our model to compute the complete reionization history, we utilize comoving simulation boxes spanning from $z=15$ to $z=5$, with an interval of $\Delta z=0.1$. However, for the purpose of this study, we essentially focus on the redshift range $z=5-6.1$.
@@
Our model utilizes the photon-conserving algorithm to construct the reionization topology within the simulation box. The ionization field relies on the ionization efficiency parameter $\zeta(M_{h},z)$, which estimates the available ionizing photons per hydrogen atom. This parameter may vary based on halo mass ($M_{h}$) and redshift ($z$). To address inhomogeneous recombinations, we adjust the ionization criteria to compensate for excess neutral atoms, while introducing small-scale fluctuations through a globally averaged clumping factor, $C_{\mathrm{HII}}$. This parameter remains highly uncertain due to the absence of direct observational measurements. Recently, a claim of relatively high clumping factor ($\sim 12$) has been made via an analytical argument applied to photoionization rate and mean free path measurements (Davies et al., 2024a), although most simulation studies suggest a value around $2-5$ (e.g. D’Aloisio et al., 2020; Asthana et al., 2025). In this study, we consistently fix this parameter at a value of 3. To calculate the recombination number density, it is necessary to consider the thermal history of the medium. Therefore, alongside tracking the ionization process, we also simulate the temperature evolution within each grid cell of our computational volume. The code is designed to automatically factor in the effects of spatially inhomogeneous reionization on temperature evolution. Specifically, when a region becomes ionized for the first time, its temperature is assumed to increase by a specific amount, denoted as $T_{\mathrm{re}}$ (Hui & Gnedin, 1997; Furlanetto & Oh, 2009; Keating et al., 2018; Maity & Choudhury, 2022a).
@@
Additionally, our model accounts for radiative feedback, which inhibits the production of ionizing photons in halos that experience gas heating. We follow the ‘step feedback’ framework introduced in Maity & Choudhury (2022a), where halos below a certain mass threshold, defined as $M_{\mathrm{min}}=\mathrm{Max}\left[M_{\mathrm{cool}},M_{J}\right]$, are assumed to retain no gas, while those above retain all their gas. Here, $M_{\mathrm{cool}}$ is the atomic cooling threshold mass, and $M_{J}$ is the Jeans mass at virial overdensity. This simplified step-function approach to radiative feedback is both efficient and effective for modeling reionization. Importantly, $M_{\mathrm{min}}$ varies with redshift and across spatial locations, especially in ionized zones where temperature-driven feedback becomes significant. The Jeans mass, being temperature-dependent, contributes to this variability. In contrast, the neutral regions are not impacted by feedback, where $M_{\mathrm{min}}$ aligns with $M_{\mathrm{cool}}$, which is roughly $\sim 3\times 10^{7}M_{\odot}$ at $z=20$ and $\sim 2\times 10^{8}M_{\odot}$ at $z=5$. In feedback-affected zones, $M_{\mathrm{min}}$ typically surpasses $10^{9}M_{\odot}$. .
@@ 2.3.3 Model for UVB Fluctuations
The fluctuations in HI photoionization rate are a crucial component in determining the Lyman-$\alpha$ optical depth during the later stages of reionization era. These mainly arise due to the discrete nature and clustering of ionizing sources which remains even after ionization equilibrium is established in post-ionization phase (Mesinger & Furlanetto, 2009). However, this is hard to accurately model within the standard semi numerical reionization framework due to the lack of small-scale information.
In this work, we deployed the model developed by Davies & Furlanetto (2016) which incorporates the effect of fluctuating mean free path on the photoionization rate. The photoionization rate in each cell ($i$) is computed by integrating the product of specific ionizing intensity (including local, $J_{\nu,i}^{\mathrm{L}}$ and non local, $J_{\nu,i}^{\mathrm{NL}}$ contribution) and interaction cross section ($\sigma_{\mathrm{HI}}$) across UVB photon frequency ($\nu$), i.e.
@@
This integration over the frequency makes the computation slightly inefficient, although it has been checked that a single frequency in the middle provides a good approximation which is sufficient for this study. To compute this relation, we need to estimate the non local contribution which is given as
@@
where $r_{ij}$ is the distance between $i$-th and $j$-th cells; $\lambda(x)$ is the mean free path at each position along the sightlines and $L_{j,\nu}$ is the specific luminosity on the corresponding ($j$-th) cell. This requires the mean free path variation for each cell that is computed as
$\lambda_{\nu}(x)\propto\lambda_{0}\leavevmode\nobreak\ \Delta^{-1}(x)\leavevmode\nobreak\ \Gamma_{\mathrm{HI}}^{2/3}(x)\leavevmode\nobreak\ \nu^{0.9}$ (7)
Here, $\lambda_{0}$ gives a proxy of averaged mean free path which can be treated as a free parameter in the model and $\Delta$ is the overdensity distribution. The dependencies on the different physical parameters via the above mentioned power law indices come from realistic assumptions about absorbing gas density distribution, however, small variations in these indices do not change the fluctuations significantly (Gaikwad et al., 2023). On the other hand, the local contribution takes into account the intensity budget within a cell which is estimated using an approximation to get a resolution independent average photoionization rate, i.e.
@@
with $\Delta x$ is the resolution of the cell and $\epsilon_{\nu,i}$ is the specific emissivity. The equation 5 and 6 can be solved iteratively in each cell to get the UVB fluctuations field.
@@ 2.3.4 Comparison with hydro skewers
In the middle and right panels of Figure 4, we show the optical depth distribution ($\tau_{i}$) at pixel scale ($4h^{-1}\mathrm{cMpc}$) and effective optical depth ($\tau_{\mathrm{eff}}$, averaged at $\Delta z=0.1$) respectively at a fiducial redshift, $z=5.0$. The distribution of opacities from hydro skewers is shown in blue, the optical depths estimated using our calibrated empirical relation on hydro density and temperature field are shown in green while the estimates from the semi numerical methods (the fiducial model has been discussed separately in the next section) are in dashed orange. We found that the fluctuations at the pixel scale are slightly lower for the semi numerical setup than the full hydro simulation which is not very surprising due to the involvement of complex astrophysical processes in those simulations and the fact that we do not introduce the additional scatter in the modified FGPA relation. However, the $\tau_{\mathrm{eff}}$ distributions are very similar to each other providing a check on the robustness of the calibration procedure. One possible reason behind these findings may be the contribution of large-scale correlations in peculiar velocities which are present in the hydro simulation, which we neglect in the semi numerical model. The peculiar velocity component is ideally a part of FGPA relation that can suppress the scatter in optical depth while averaging along the line of sight. However, this needs to be investigated in a separate study later after a proper comparison of velocity fields between the semi numerical model and hydro simulation.
Figure 6: The snapshots of density ($\Delta$), temperature ($T$), UVB fluctuations ($\Gamma_{\mathrm{HI}}/\langle\Gamma_{\mathrm{HI}}\rangle$) and Ly-$\alpha$ optical depth ($\tau_{i}$), gradually from left to right using a fiducial setup of the calibrated semi numerical model at redshift, $z=5.4$.
@@ 3 Demonstration with Fiducial Model
In this section, we demonstrate the features of our model assuming a fiducial reionization model. We choose a fiducial model from the allowed posterior parameter spaces satisfying a variety of available observational constraints (Maity & Choudhury, 2022b) including CMB scattering optical depth, UV luminosity functions, and low density IGM temperature estimates. In Figure 5, we show the evolution of neutral fractions (reionization history), mean temperature evolution, and power law index ($\gamma$) evolution of $T-\Delta$ relation for our fiducial model. We show the recent observational constraints on these quantities as well i.e. (neutral fractions; Davies et al., 2018b; Greig et al., 2022; Jin et al., 2023; Zhu et al., 2024; Spina et al., 2024) and (IGM temperature; Gaikwad et al., 2020). It is apparent that the fiducial model obeys all the constraints reasonably well. We also show two variants of the model using different temperature increment ($T_{\mathrm{re}}$) values representing two extremes. We choose these models such that they just become consistent with the observational constraints on mean IGM temperature from the two ends. Later, we’ll use these models to marginalize over different allowed temperatures while estimating photoionization rate and mean free path.
@@
Next, we study the effects of varying the parameters which are responsible for generating the UVB fluctuations. In Figure 6, we show the snapshots of various physical quantities for the fiducial model at redshift $z=5.4$. These are density ($\Delta$), temperature ($T$), UVB fluctuations ($\Gamma_{\mathrm{HI}}/\langle\Gamma_{\mathrm{HI}}\rangle$) and the Ly-$\alpha$ optical depth ($\tau$) from left to right respectively. The correlation between different fields are quite apparent from the snapshots. The overdense regions are ionized at a relatively earlier time, allowing more time for cooling. Hence, the temperature in those regions are comparatively lower. The photoionization rate on the other hand is larger for high density regions. This is expected as high density regions naturally host galaxies and stars which are supposed to be dominant sources of ionizing photons. In this fiducial scenario, we have used $\lambda_{0}=10\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$ for the UVB field. For this small mean free path, the opacity fluctuations are mainly dominated by the ionization background fluctuation. However, the strength of this fluctuation can vary depending on the mean free path parameter which we discuss later.
@@
In Figure 7, we demonstrate the effect of UVB and temperature fluctuations on the Ly-$\alpha$ opacities. At the left panel, we show the case where no UVB or temperature fluctuations are included and the opacities are driven solely by the density field. We include temperature fluctuations in the next panel which increases the fluctuations as overdense regions become more opaque due to lower temperature and vice versa. On the third panel, we incorporate weakly fluctuating UVB field which slightly neutralises the fluctuations. This happens as the high density regions have higher photoionization rates contributing lower opacities which further results in smoothening the optical depth field. At the rightmost panel, the fluctuations are again increased due to a strong UVB field. In Figure 8, we show the cumulative probability distributions (CDF) of the effective optical depth ($\tau_{\mathrm{eff}}$) for different mean photoionization rate (in $10^{-12}/s$ i.e. $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$, at left) and effective mean free path ($\lambda_{0}$, at right). It can be seen that the mean photoionization rate mainly shifts the mean value of $\tau_{\mathrm{eff}}$ while $\lambda_{0}$ contributes to the amount of fluctuation in the distribution. A lower value of mean photoionization rate corresponds to a higher effective optical depth distribution and vice versa. This is also expected from equation 2 where these two quantities are inversely related to each other. Physically, the photoionization rate also controls the transmission flux i.e., a higher photoionization provides a higher transmission which corresponds to a lower optical depth. On the other hand, mean free path essentially provides an estimate how far a photon can travel before getting absorbed by the residual neutral gas. Hence, a higher value of $\lambda_{0}$ results in a smoother transmission having less amount of fluctuations due to the absorption which further narrows down the optical depth distribution.
@@
Figure 8: The figure shows the cumulative distribution (CDF) of $\tau_{\mathrm{eff}}$ at redshift, $z=5.4$ for a variety of parameter values. The left panel shows the variation of CDF for different mean photoionization rate (in $10^{-12}/s$ i.e. $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$). The middle panel shows the CDF variation with effective mean free path ($\lambda_{0}$ in $h^{-1}\mathrm{cMpc}$), the right panel shows the CDF variation with different reionization temperature increment ($T_{\mathrm{re}}$).
@@
Figure 9: Left Panel: An example of Probability Distribution Functions (PDF) from the data and the fiducial model at redshift, $z=5.4$. The blue line shows the histogram obtained from the data while the dashed black line shows the fit using Gaussian KDE. Similarly, the red line shows the distribution obtained from our fiducial model. For the likelihood analysis, we use 15 equally spaced bins as marked by the black solid points. Right Panel: The normalized covariance matrix at $z=5.4$.
@@
Figure 10: The plots of cumulative probability distribution of $\tau_{\mathrm{eff}}$ for different redshifts ($z=5.0-6.1$). Each panel shows the combined CDFs of the bestfit model using 500 random realizations (red) along with the individual realization (cyan) and observational data (black). We show both the optimistic and pessimistic scenarios for the data.

### v2

@@ 2.3.1 Generating density and collapsed halo field
For generating the cosmological density field, we used the approach adopted in the publicly available semi-numerical reionization model 21cmFAST (Mesinger et al., 2011). The initial conditions on density and velocity fields are generated in Lagrangian space following the gravitational evolution of the fields via Zel’Dovich Approximation (Zel’dovich, 1970). This approximation has been proven to be successful in closely resembling the full N-body simulations with less computational cost. We used a simulation of box size $256\penalty 10000\ h^{-1}\mathrm{cMpc}$, which is sufficiently large to serve the purpose of this study. After having the particle fields, we gridded those at a high resolution ($0.5h^{-1}\mathrm{cMpc}$) and then averaged them to $\Delta x=4h^{-1}\mathrm{cMpc}$, which is the resolution we adopted throughout the study.
@@
Besides, to construct the collapsed halo mass fraction, we adopted the Excursion Set Formalism (ESF-L) in Lagrangian space, following another semi-numerical reionization model AMBER (Trac et al., 2022). To be specific, the formalism uses Extended Press Schecter theory, which computes the collapsed mass fraction ($f_{\mathrm{coll}}$) above a minimum threshold halo mass ($M_{\mathrm{min}}$) given the filtered overdensity at a cosmological scale. Here, we used a spherical tophat filter for filtering the overdensity at a certain scale ($2\penalty 10000\ \mathrm{cMpc}$), which provides a good match with the full N-body simulation. At each redshift, we generated a collapsed fraction field for a logarithmic minimum mass ($\log M_{\mathrm{min}}$) interval of $0.1$ in the range 7127-12.
@@
In the left panel of Figure 4, we show the overdensity distribution of the Nyx simulation (blue) and the generated overdensities using semi-numerical method (dashed orange) as discussed above. We checked that the natures of these two distributions are in agreement with each other. This resemblance is useful to avoid any discrepancies in optical depth due to the background density field.
2.3.2 Semi-numerical reionization model
@@
We utilized an explicitly photon-conserving semi-numerical reionization model, SCRIPT, which provides the ionization and thermal state of the universe in a cosmologically representative simulation volume (Choudhury & Paranjape, 2018; Maity & Choudhury, 2022a). To initiate the model, we provided inputs comprising the density field and the distribution of collapsed halos capable of emitting ionizing radiation as generated by the method discussed earlier. For our model to compute the complete reionization history, we utilized comoving simulation boxes spanning from $z=15$ to $z=5$, with an interval of $\Delta z=0.1$. However, for the purpose of this study, we essentially focused on the redshift range z=56.1z=5-6.1.
@@
Our model utilizes the photon-conserving algorithm to construct the reionization topology within the simulation box. The ionization field relies on the ionization efficiency parameter $\zeta(M_{h},z)$, which estimates the available ionizing photons per hydrogen atom. This parameter may vary based on halo mass ($M_{h}$) and redshift ($z$). To address inhomogeneous recombinations, we adjust the ionization criteria to compensate for excess neutral atoms, while introducing small-scale fluctuations through a globally averaged clumping factor, $C_{\mathrm{HII}}$. This parameter remains highly uncertain due to the absence of direct observational measurements. Recently, a claim of a relatively high clumping factor ($\sim 12$) has been made via an analytical argument applied to photoionization rate and mean free path measurements (Davies et al., 2024a), although most simulation studies suggest a value around 252-5 (e.g. D’Aloisio et al., 2020; Asthana et al., 2025). In this study, we consistently fixed this parameter at a value of 3. To calculate the recombination number density, it is necessary to consider the thermal history of the medium. Therefore, alongside tracking the ionization process, we also simulated the temperature evolution within each grid cell of our computational volume. The code is designed to automatically factor in the effects of spatially inhomogeneous reionization on temperature evolution. Specifically, when a region becomes ionized for the first time, its temperature is assumed to increase by a specific amount, denoted as $T_{\mathrm{re}}$ (Hui & Gnedin, 1997; Furlanetto & Oh, 2009; Keating et al., 2018; Maity & Choudhury, 2022a).
@@
Additionally, our model accounts for radiative feedback, which inhibits the production of ionizing photons in halos that experience gas heating. We followed the ‘step feedback’ framework introduced in Maity & Choudhury (2022a), where halos below a certain mass threshold, defined as $M_{\mathrm{min}}=\mathrm{Max}\left[M_{\mathrm{cool}},M_{J}\right]$, are assumed to retain no gas, while those above retain all their gas. Here, $M_{\mathrm{cool}}$ is the atomic cooling threshold mass, and $M_{J}$ is the Jeans mass at virial overdensity. This simplified step-function approach to radiative feedback is both efficient and effective for modeling reionization. Importantly, $M_{\mathrm{min}}$ varies with redshift and across spatial locations, especially in ionized zones where temperature-driven feedback becomes significant. The Jeans mass, being temperature-dependent, contributes to this variability. In contrast, the neutral regions are not impacted by feedback, where $M_{\mathrm{min}}$ aligns with $M_{\mathrm{cool}}$, which is roughly $\sim 3\times 10^{7}M_{\odot}$ at $z=20$ and $\sim 2\times 10^{8}M_{\odot}$ at $z=5$. In feedback-affected zones, $M_{\mathrm{min}}$ typically surpasses $10^{9}M_{\odot}$. .
@@ 2.3.3 Model for UVB fluctuations
The fluctuations in HI photoionization rate are a crucial component in determining the Lyman-$\alpha$ optical depth during the later stages of the reionization era. These mainly arise due to the discrete nature and clustering of ionizing sources, which remains even after ionization equilibrium is established in the post-ionization phase (Mesinger & Furlanetto, 2009). However, this is hard to accurately model within the standard semi-numerical reionization framework due to the lack of small-scale information.
In this work, we deployed the model developed by Davies & Furlanetto (2016), which incorporates the effect of fluctuating mean free path on the photoionization rate. The photoionization rate in each cell ($i$) is computed by integrating the product of specific ionizing intensity (including local, $J_{\nu,i}^{\mathrm{L}}$ and non local, $J_{\nu,i}^{\mathrm{NL}}$ contribution) and interaction cross section ($\sigma_{\mathrm{HI}}$) across UVB photon frequency ($\nu$), i.e.
@@
This integration over the frequency makes the computation slightly inefficient, although it has been checked that a single frequency ($\nu\sim 1.32\nu_{\mathrm{HI}}$) in the middle provides a good approximation (Davies & Furlanetto, 2016), which is sufficient for this study. To compute this relation, we need to estimate the non-local contribution, which is given as
@@
where $r_{ij}$ is the distance between $i$-th and $j$-th cells; $\lambda(x)$ is the mean free path at each position along the sightlines, and $L_{j,\nu}$ is the specific luminosity on the corresponding ($j$-th) cell. This requires the mean free path variation for each cell that is computed as
$\lambda_{\nu}(x)\propto\lambda_{0}\penalty 10000\ \Delta^{-1}(x)\penalty 10000\ \Gamma_{\mathrm{HI}}^{2/3}(x)\penalty 10000\ \nu^{0.9}$ (7)
Here, $\lambda_{0}$ gives a proxy of the averaged mean free path, which can be treated as a free parameter in the model, and $\Delta$ is the overdensity distribution. The dependencies on the different physical parameters via the above-mentioned power law indices come from realistic assumptions about absorbing gas density distribution; however, small variations in these indices do not change the fluctuations significantly (Gaikwad et al., 2023). On the other hand, the local contribution takes into account the intensity budget within a cell, which is estimated using an approximation (Davies & Furlanetto, 2016) to get a resolution-independent average photoionization rate, i.e.
@@
with $\Delta x$ is the resolution of the cell and $\epsilon_{\nu,i}$ is the specific emissivity. The equations 5 and 6 can be solved iteratively in each cell to get the UVB fluctuations field.
@@ 2.3.4 Comparison with hydro skewers
In the middle and right panels of Figure 4, we show the optical depth distribution ($\tau_{i}$) at pixel scale ($4h^{-1}\mathrm{cMpc}$) and effective optical depth ($\tau_{\mathrm{eff}}$, averaged at $\Delta z=0.1$) respectively at a fiducial redshift, $z=5.0$. The distribution of opacities from hydro skewers is shown in blue, the optical depths estimated using our calibrated empirical relation on hydro density and temperature field are shown in green, while the estimates from the semi-numerical methods (the fiducial model has been discussed separately in the next section) are in dashed orange. We found that the fluctuations at the pixel scale are slightly lower for the semi-numerical setup than the full hydro simulation, which is not very surprising due to the involvement of complex astrophysical processes in those simulations and the fact that we do not introduce the additional scatter in the modified FGPA relation. However, the $\tau_{\mathrm{eff}}$ distributions are very similar to each other, providing a check on the robustness of the calibration procedure. One possible reason behind these findings may be the contribution of large-scale correlations in peculiar velocities, which are present in the hydro simulation, which we neglect in the semi-numerical model. The peculiar velocity component is ideally a part of the FGPA relation that can suppress the scatter in optical depth while averaging along the line of sight. However, this needs to be investigated in a separate study later, after a proper comparison of velocity fields between the semi-numerical model and hydro simulation.
Figure 6: The snapshots of density ($\Delta$), temperature ($T$), UVB fluctuations ($\Gamma_{\mathrm{HI}}/\langle\Gamma_{\mathrm{HI}}\rangle$) and Ly-$\alpha$ optical depth ($\tau_{i}$), gradually from left to right using a fiducial setup of the calibrated semi-numerical model at redshift, $z=5.4$.
@@ 3 Demonstration with fiducial model
In this section, we demonstrate the features of our model assuming a fiducial reionization model. We chose a fiducial model from the allowed posterior parameter spaces satisfying a variety of available observational constraints (Maity & Choudhury, 2022b), including CMB scattering optical depth, UV luminosity functions, and low-density IGM temperature estimates. In Figure 5, we show the evolution of neutral fractions (reionization history), mean temperature evolution, and power law index ($\gamma$) evolution of $T-\Delta$ relation for our fiducial model. We show the recent observational constraints on these quantities as well, i.e. (neutral fractions; Davies et al., 2018b; Greig et al., 2022; Jin et al., 2023; Zhu et al., 2024; Spina et al., 2024) and (IGM temperature; Gaikwad et al., 2020). It is apparent that the fiducial model obeys all the constraints reasonably well. We also show two variants of the model using different temperature increments ($T_{\mathrm{re}}$) values representing two extremes. We chose these models such that they just become consistent with the observational constraints on mean IGM temperature from the two ends. Later, we’ll use these models to marginalize over different allowed temperatures while estimating the photoionization rate and mean free path.
@@
Next, we studied the effects of varying the parameters that are responsible for generating the UVB fluctuations. In Figure 6, we show the snapshots of various physical quantities for the fiducial model at redshift $z=5.4$. These are density ($\Delta$), temperature ($T$), UVB fluctuations ($\Gamma_{\mathrm{HI}}/\langle\Gamma_{\mathrm{HI}}\rangle$) and the Ly-$\alpha$ optical depth ($\tau$) from left to right respectively. The correlation between different fields is quite apparent from the snapshots. The overdense regions are ionized at a relatively earlier time, allowing more time for cooling. Hence, the temperatures in those regions are comparatively lower. The photoionization rate, on the other hand, is larger for high density regions. This is expected as high-density regions naturally host galaxies and stars, which are supposed to be dominant sources of ionizing photons. In this fiducial scenario, we have used $\lambda_{0}=10\penalty 10000\ h^{-1}\mathrm{cMpc}$ for the UVB field. For this small mean free path, the opacity fluctuations are mainly dominated by the ionization background fluctuation. However, the strength of this fluctuation can vary depending on the mean free path parameter, which we discuss later.
@@
In Figure 7, we demonstrated the effect of UVB and temperature fluctuations on the Ly-$\alpha$ opacities. At the left panel, we show the case where no UVB or temperature fluctuations are included, and the opacities are driven solely by the density field. We included temperature fluctuations in the next panel, which increases the fluctuations as overdense regions become more opaque due to lower temperature and vice versa. On the third panel, we incorporated a weakly fluctuating UVB field, which slightly neutralises the fluctuations. This happens as the high-density regions have higher photoionization rates, contributing to lower opacities, which further results in smoothening the optical depth field. At the rightmost panel, the fluctuations are again increased due to a strong UVB field. In Figure 8, we show the cumulative probability distributions (CDF) of the effective optical depth ($\tau_{\mathrm{eff}}$) for different mean photoionization rate (in $10^{-12}/s$ i.e. $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$, at left) and effective mean free path ($\lambda_{0}$, at right). It can be seen that the mean photoionization rate mainly shifts the mean value of $\tau_{\mathrm{eff}}$ while $\lambda_{0}$ contributes to the amount of fluctuation in the distribution. A lower value of the mean photoionization rate corresponds to a higher effective optical depth distribution and vice versa. This is also expected from equation 2, where these two quantities are inversely related to each other. Physically, the photoionization rate also controls the transmission flux, i.e., a higher photoionization rate provides a higher transmission, which corresponds to a lower optical depth. On the other hand, the mean free path essentially provides an estimate of how far a photon can travel before getting absorbed by the residual neutral gas. Hence, a higher value of $\lambda_{0}$ results in a smoother transmission having less amount of fluctuations due to the absorption, which further narrows down the optical depth distribution.
@@
Figure 8: The figure shows the cumulative distribution (CDF) of $\tau_{\mathrm{eff}}$ at redshift, $z=5.4$ for a variety of parameter values. The left panel shows the variation of CDF for different mean photoionization rate (in $10^{-12}/s$, i.e., $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$). The middle panel shows the CDF variation with effective mean free path ($\lambda_{0}$ in $h^{-1}\mathrm{cMpc}$), the right panel shows the CDF variation with different reionization temperature increment ($T_{\mathrm{re}}$).
@@
Figure 9: Left Panel: An example of Probability Distribution Functions (PDF) from the data and the fiducial model at redshift, $z=5.4$. The blue line shows the histogram obtained from the data, while the dashed black line shows the fit using Gaussian KDE. Similarly, the red line shows the distribution obtained from our fiducial model. For the likelihood analysis, we use 15 equally spaced bins as marked by the black solid points. Right Panel: The normalized covariance matrix at $z=5.4$.
@@
Figure 10: The CDF plots of $\tau_{\mathrm{eff}}$ for different redshifts (z=5.06.1z=5.0-6.1). Each panel shows the combined CDFs of the bestfit model using 500 random realizations (red) along with the individual realization (cyan) and observational data (black). We show both the optimistic and pessimistic scenarios for the data.